r+t/2=7 ;t
Twice the sum of 3 consecutive number odd integers is 150. Find the 3 integers.
I know you start off the problem like this.
x+x+2+x+4=150
then I get 3x+6=150
then 3x=154
... I'm not sure where I went or if I went wrong.

Found 2 solutions by scott8148, Alan3354:
Answer by scott8148(6628)

I don't see a TWICE anywhere...

Answer by Alan3354(69443)

r+t/2=7 ;t
-------
If you mean solve for t
2r + t = 14
t = 14-2r
-------------
Twice the sum of 3 consecutive number odd integers is 150. Find the 3 integers.
I know you start off the problem like this.
x+x+2+x+4=150
Close, but it says "Twice", so it's
2*(x+x+2+x+4)=150
-------------------
then I get 3x+6=150
---------------
It's twice that, 6x+12=150
then 3x=154 (should be 144, but that's irrelevant)
... I'm not sure where I went or if I went wrong
-------------
6x+12=150
6x = 138
x = 23
So the integers are 23,25,27
They add up to 75, and twice that is 150.